Differential Equation

Please answer both of the questions below Thanks!

Solve the given initial value problem. y + 36y = 0; y(0) = 3, y(O) = 5 x(t) =

Find a general solution to the differential equation using the method of variation of parameters. y + 2y +y=2e -t The gene
Answer #1

= o y +36y=0 Y+36y=0 y 10) = 410) 25 Ausiliary Equation m²+36=0 m² = -36 m= +61 solution will be yt) = C, coset & Casinot gIt y = pot Amiliary Can let y+ 2ylty = xit m² tomt 1 - 0 (m+ = m=- -| Saltion will be complementry function, C.F. = (atletetand u= (- 4 gt) at yiya - Ye Yi - tot eetat et [- tette t] + te txost (tzt dt - text, p2t + toat v = 1 = Auto = 5-34 (-atest